### v1

# The Second Machine Turn: From Checking Proofs to Creating Concepts

The Second Machine Turn: From Checking Proofs to Creating Concepts
Asvin G. [^1] The author thanks Claude Opus 4 (Anthropic) for extensive discussions that helped develop and refine these ideas.
In 1854 George Boole extended Aristotle’s syllogisms with an algebra of equations representing logical thought, showing that reasoning itself could be reduced to rule-bound calculation. David Hilbert, sensing the power of those rules, called for an engine that would mechanize certainty. Kurt Gödel discovered the impossibility of such certainty, yet Alan Turing built the universal engine anyway: a machine that could, in principle, instantiate any formal system. Today the descendants of that engine — Lean, Coq, Isabelle — are moving from specialist enclaves toward everyday use, promising a future in which every inference can be mechanically checked. But a deeper question follows. Once the mechanism is reliable, who decides where to turn it? Can we mechanize insight?
This is not idle speculation. As I write this in mid-2025, AI systems are solving both International Mathematical Olympiad problems and graduate level mathematics while the community is simultaneously formalizing vast swaths of existing mathematics. Given the rate of progress in the field, is it possible that within a year or two, these systems may match or exceed human mathematicians at the full spectrum of mathematical work: not just calculation or verification, but the creative acts of conjecture, proof, and theory-building that have always seemed uniquely human? The answer is far from obvious and where these questions deserve our utmost consideration.
If we are to be active participants in the next revolution rather than mere bystanders, mathematicians must now do for the idea of “concept”; the units in which insights are measured, what Euclid once did for proof: give it a precise, mathematical formulation. Defining such a grammar - how concepts are discovered, compressed, ported, and retired - is not a task for engineers alone. It is a mathematical problem whose solution will determine whether future intelligences reason with us or past us. This essay takes a first step towards this goal. But to understand where we are headed, we must first understand where we have been. The story of mathematics has always been a story of co-evolution with our tools.
The Shifting Landscape of Mathematical Labor
Before the nineteenth century, calculation was the essence of mathematics. Gauss spent years computing the orbits of asteroids. Euler calculated decimal expansions of $\pi$ by hand. When Legendre wanted to understand the distribution of primes, he manually counted them up to 400,000, then extended Gauss’s counts to a million. These weren’t preliminary exercises—they were the main work. Mathematical reputation rested as much on computational stamina as theoretical insight.
The electronic computer changed everything. Suddenly, any graduate student could exceed Gauss’s computational output in an afternoon. The mathematical community’s response was revealing: we collectively shifted our valuation of mathematical labor. The focus moved from computation to proof, particularly to abstract, conceptual arguments that seemed furthest from mechanical calculation. The Bourbaki movement, with its emphasis on axiomatic foundations and structural clarity, represented one pole of this shift. Mathematics, in this view, was about pure logical relationships and conceptual architecture, not computational grunt work.
Yet this shift created its own problems. As mathematical arguments grew more abstract and intricate, verification became increasingly difficult. The classification of finite simple groups required tens of thousands of journal pages. Who could check such a proof? Who could even read it? The four-color theorem’s computer-assisted proof in 1976 brought the crisis into sharp focus: was a proof still a proof if no human could verify it directly?
The response, still ongoing, has been a gradual turn toward formalization. If humans struggle to verify complex proofs, perhaps machines can help—not by finding proofs, but by checking them. Projects like Lean, Coq, and Isabelle promise a future where mathematical arguments are not just correct but verifiably correct, checkable down to the last logical step.
Now we stand at another precipice. What happens when machines don’t just check proofs but create them?
Mathematics as Experimental Science
To answer this question, we need to reconsider what mathematicians actually do. In recent work [g2025unreasonableeffectivenessmathematicalexperiments], we’ve argued that mathematics is best understood not as the discovery of eternal truths but as an experimental science exploring patterns in computation.
The core insight is this: Science is the process of discovering interesting artifacts and then explaining them. Physicists discover artifacts in the physical world—the perihelion of Mercury, the spectrum of hydrogen, the cosmic microwave background—and build theories to explain them. Mathematicians discover artifacts in the computational universe—patterns that emerge when symbols are manipulated according to rules—and build theories to explain these patterns.
Consider Euler’s famous claim that $1+2+3+\cdots=-1/12$. This wasn’t mysticism or error; it was an experimental observation. Euler had discovered that treating divergent series as if they converged led to consistent, useful results across multiple problems. The artifact was this surprising regularity. The explanation would come a century later with the theory of analytic continuation.
From this perspective, mathematical progress follows a natural cycle: compute and explore, discovering unexpected patterns; develop concepts and frameworks to explain these patterns; use proof to verify that theories correctly predict the observed phenomena. This framework dissolves many traditional puzzles. Why is mathematics so effective in physics? Because physics, too, relies fundamentally on computation. Why do some mathematical concepts prove so broadly applicable? Because they capture computational patterns that appear across many contexts. Mathematical truth isn’t about correspondence to platonic reality but about successful prediction of computational outcomes.
Most importantly for our current moment: if mathematics is fundamentally about recognizing patterns in computation and building frameworks to explain them, then the question isn’t whether AI can do mathematics, but what kind of mathematics it will do—and crucially, what role remains for humans in this new landscape.
How Machines Might Learn Math
To understand where AI might take mathematics, we must first understand how these systems work. The breakthrough came not from mimicking human logical reasoning but from an insight about learning: the power of self-play and exploration.
In 2016, DeepMind’s AlphaGo [silver2017mastering] defeated Lee Sedol, one of the world’s strongest Go players. The victory shocked experts who had predicted such an achievement was decades away. Go’s branching factor—the number of possible moves at each turn—is around 250, making brute-force search impossible even with universe-scale computing. How did AlphaGo succeed where traditional approaches failed?
The answer lies in combining two mechanisms. First, a neural network learns a policy function: a probability distribution over plausible moves in any given position. Through playing countless games against itself, this network develops what we might call intuition—strategic principles that transcend individual positions. Second, when facing a specific position, the system searches through possible continuations, using its learned policy to guide exploration toward the most promising branches. The policy function captures general principles while the search localizes these principles to the specific position at hand. In effect, search functions as a computational process interpretable by the policy function, allowing the system to understand local structure through the lens of learned concepts.
This is not mere anthropomorphism. Studies of AlphaZero [mcgrath2022acquisition, schut2025bridging] reveal that these systems develop internal representations corresponding to human chess concepts: ”good-and-bad-bishops,” ”mate-threats,” ”open-files.” The neural networks aren’t just memorizing positions—they’re discovering the same abstract patterns that human masters use to organize their understanding of the game.
Modern large language models have followed a strikingly parallel path. During training, these models develop their own policy function: a probability distribution over likely next words given a piece of text. Recent developments in reasoning-enabled models like o1 have introduced explicit search into language generation. When solving a problem, these models don’t simply output the first plausible answer. Instead, they explore multiple reasoning paths, evaluate their promise, and navigate toward solutions much as AlphaZero navigates toward winning positions. Research [templeton2024scaling, ameisen2025circuit] shows that LLMs similarly learn interpretable concepts: ”Golden-Gate-Bridge,” ”capital-city,” and countless others emerge naturally from the training process.
Mathematics presents a natural frontier for these techniques. Two major efforts suggest how AI might tackle mathematical reasoning, each with distinct advantages and limitations.
The first approach, exemplified by DeepMind’s AlphaProof, adapts the AlphaGo architecture directly to theorem proving. Working within formal systems like Lean, the AI attempts to prove theorems by searching through the space of valid logical moves. The policy network learns which proof steps are promising; the search component verifies that these steps actually work. Initial training comes from a corpus of human-formalized proofs, but the system improves through self-play: attempting proofs, learning from successes and failures.
This approach offers the overwhelming advantage of certainty. Every proof the system produces is mechanically checkable, eliminating concerns about subtle errors or hallucinated arguments. This certainty enables self-play: the system can generate its own training data by attempting new proofs and learning from the results. Early results have been impressive: AlphaProof has solved International Mathematical Olympiad problems that challenge the world’s brightest young mathematicians.
Yet formalization creates its own challenges. Most human mathematics isn’t written in formal languages like Lean but in a rich mixture of natural language, symbolic notation, and implicit shared understanding. The gap between $\forall\epsilon>0,\exists\delta>0$ and ”for any desired accuracy, we can find a sufficiently close…” encompasses more than notation. Converting between these representations requires sophisticated translation: ”autoformalization” in one direction; making formal proofs human-readable in the other. Scaling either task remains a significant challenge.
The second approach embraces natural language directly. Recent models demonstrate graduate-level performance on mathematical competitions [EpochLLMBenchmarkingHub2024], generating proofs in the same mixture of formal notation and explanatory prose that mathematicians use. These systems can explain their reasoning, making their proofs more accessible and their errors more diagnosable. Training data abounds in textbooks, papers, and online resources.
However, verification becomes the central challenge. Without formal proof checking, we lose the clean signal that enables self-play. The model might generate plausible-sounding but subtly flawed arguments. Self-improvement requires external validation or sophisticated methods for assessing proof correctness in natural language, a problem that remains largely unsolved.
Both approaches likely build genuine mathematical understanding. The evidence from game-playing systems and language models strongly suggests that sufficiently powerful neural networks trained on mathematical tasks will develop internal representations of mathematical concepts—not just syntactic manipulation rules but semantic understanding of continuity, symmetry, dimensionality, and other core ideas. Just as AlphaZero discovers open files and pins, a mathematical AI might implicitly grasp that compactness and finiteness share deep structural similarities — ”compactness as a way of controlling infinite process through finitary means”. In topology, this manifests as finite subcovers; in algebra, as finite generators; in algebraic geometry, as properness. The AI might cluster these concepts in its representational space without explicitly naming the pattern.
Yet mathematics poses unique challenges beyond those faced in game-playing. The search space explodes exponentially with proof length, far more severely than in board games where positions have bounded complexity. A chess game might last 40 moves with perhaps 30 legal moves per position. A significant mathematical proof might require hundreds of steps, each chosen from an effectively infinite set of possibilities. Moreover, mathematical proofs demand perfect correctness throughout. In Go, a suboptimal move might reduce winning probability but rarely proves fatal. In mathematics, a single flawed inference invalidates everything that follows.
This exponential search difficulty raises a profound question: How do human mathematicians navigate this vast space successfully? The combination of learned intuition and systematic search seems insufficient to explain how mathematicians discover proofs that sometimes span hundreds of pages, maintaining logical coherence throughout while exploring genuinely novel territory.
The answer, as we’ll see, lies in the same limitation that makes brute-force proof search impossible: the necessity of conceptual compression.
Reshaping the Search Landscape
Open any mathematics textbook and you’ll find definitions, theorems, and proofs marching in orderly succession. This formal presentation is the logical skeleton of mathematics—necessary for rigor, but as far from the living discipline as a skeleton is from a dancer. When we learn mathematics, we don’t absorb these logical structures directly. Instead, we use them to bootstrap conceptual understanding, running mental experiments, building intuitions, making connections that exist nowhere on the page.
The same is true in reverse. When mathematicians create new mathematics, they don’t proceed by systematic logical search. The reason is computational impossibility. Even a modest theorem might require a proof of twenty steps, each chosen from hundreds of possible moves. The search space explodes exponentially: a proof with merely $20$ lines could easily have at least $100^{20}$ possibilities, more than the number of atoms in the observable universe. No amount of computational power could explore such spaces by brute force.
How then do mathematicians find proofs? Through the same mechanism that makes mathematics learnable: conceptual compression. Consider how the notion of integration transforms calculation. Computing the area under a curve through Riemann sums requires evaluating an infinite summation and then taking a limit — often a computational nightmare even in simple cases. But recognizing that integration connects to antiderivatives through the Fundamental Theorem of Calculus transforms the problem entirely. What was an infinite process becomes a finite calculation: evaluate the antiderivative at two points and subtract. The concept of integration doesn’t just organize our thinking; it narrows down a potentially unbounded computation into an almost constant calculation.
The entire edifice of homological algebra arose from similar computational pressures. Early algebraic topologists found themselves drowning in elaborate calculations with chain complexes, each proof requiring meticulous bookkeeping. The introduction of derived categories and spectral sequences didn’t just organize these calculations—it often eliminated them entirely. Grothendieck’s general notion of cohomology as a derived functor simplified the comparison of de-Rham cohomology with singular cohomology from a complicated calculation into an almost inevitable consequence of a few general principles.
Each definition creates a new conceptual layer, compressing complex patterns into manipulable units. Each lemma builds a reusable component. The impossible search through $100^{20}$ possibilities becomes a guided journey through perhaps ten conceptual moves, each itself requiring only local exploration. A proof isn’t a path through logical space but a network of conceptual compressions, each node reducing local complexity to manageable proportions.
Andrew Wiles’s proof of Fermat’s Last Theorem illustrates this process at its most sophisticated. Wiles didn’t attack the problem through systematic search. Instead, he recognized that the Taniyama-Shimura conjecture might provide the conceptual bridge needed, itself a compression linking elliptic curves to modular forms. For seven years, he explored this conceptual landscape, developing new tools when existing ones proved insufficient. When a subtle gap emerged in 1993, the conceptual scaffold remained sound. He knew where he was trying to go; he needed only to find a different path through the compressed landscape. Analyzing precisely why his proof failed in conceptual terms, he developed a new technique — a different instantiation of his original strategy — and completed the proof in 1995.
This resilience to local errors reveals why explicit concepts matter. With only implicit understanding, a flaw anywhere might require starting over. But conceptual compression creates modular structures: fix one component without rebuilding the entire edifice.
Concepts are not luxuries that make mathematics prettier or easier to understand. They are computational necessities, the only way minds, human or artificial, can navigate the exponential explosion of formal possibilities. Without the ability to recognize patterns, create abstractions, and work at multiple levels simultaneously, mathematical progress would be impossible.
This is why proofs that emerge from pure computational search often feel unsatisfying. The computer-assisted proof of the four-color theorem verified the result but provided no conceptual understanding, no compression of the search space that human minds can grasp and reuse. Such proofs are logic without understanding, verification without insight.
The Necessity of Names
Chess and Go are games with fixed rules and stable conceptual vocabularies. While each position presents novel tactical challenges, the underlying patterns—pins, forks, sacrifices—remain constant across millennia. Concepts emerge implicitly through play and need never be made explicit; the game itself provides complete feedback.
Mathematics operates in a fundamentally different regime. Each new domain brings not just new problems but new modes of reasoning entirely. The introduction of probabilistic methods into combinatorics by Erdős wasn’t a recombination of existing tactics but the importation of an alien conceptual framework that transformed the field. When algebraic geometers began applying their methods to matroids, they weren’t instantiating familiar patterns in new positions but discovering that abstract structural similarities admitted powerful new tools. Mathematics continuously develops new forms of conceptual compression, each opening previously unimaginable approaches.
In practice, mathematicians develop understanding by considering simple cases, observing patterns in calculations, and proving lemmas that validate their intuitions. Current AI systems show promise in automating this exploration, particularly in formal systems like Lean where verification is guaranteed. The principles of mathematical exploration—vary parameters, examine edge cases, seek connections—appear sufficiently universal that machine learning could capture them.
But a crucial challenge remains: extracting useful concepts from such exploration. An AI system might implicitly recognize that compactness and properness share deep structural features with finiteness: both involve controlling infinite processes through finite data. This implicit clustering in representational space captures a genuine mathematical insight. Yet without crystallizing this pattern into explicit definitions and theorems (a compact discrete space is finite, proper morphisms have finite dimensional cohomology), the insight remains locked within the specific contexts where it was learned.
Consider how Grothendieck revolutionized algebraic geometry. The Italian geometers had brilliant geometric intuitions but lacked precise language to articulate them. Grothendieck’s schemes seemed absurdly abstract to contemporaries, yet by making explicit what had been implicit, by providing precise definitions for patterns others had glimpsed but couldn’t articulate, he created conceptual technology that offered the porting of traditional notions (a topology for real manifolds) to completely novel frameworks (a Grothendieck topology for varieties over finite fields, an apparently completely discrete setting) by the simple observation that an open set of a topological space need not be a subset. The explicitness framework here offered both the scope for iterative improvement as well as absorption by a broader community that eventually allowed for the proof of the Weil Conjectures through the joint effort of an entire community of mathematicians.
This is what explicit concepts provide: not just names but handles for manipulation, interfaces for composition. A neural network might learn to apply continuity in a thousand specific cases. But only by crystallizing this pattern into an explicit definition—”for every $\epsilon>0$, there exists $\delta>0$…”—does it become a tool others can grasp, combine with other concepts, and use to explore entirely new territories.
The limitation of implicit representations runs deeper than communication. Neural networks learn patterns in ways inherently tied to their training distribution. An AI trained on topological proofs might develop implicit representations of compactness perfectly suited to that domain. But these representations encode not just the abstract concept but also the specific contexts, notations, and proof techniques of topology. Extracting the pure pattern—the one that would generalize to compactness in logic, analysis, or algebraic geometry—requires disentangling concept from context, signal from noise.
Moreover, explicit concepts enable the modular error correction that makes mathematical progress possible. When Wiles found a gap in his proof, he could diagnose precisely which component had failed because his conceptual structure was explicit. With only implicit understanding distributed across network weights, how would an AI identify and repair flawed reasoning? The very features that make neural representations powerful—their distributed, context-sensitive nature—make them fragile and opaque.
The gap between pattern recognition and concept formation is ancient. Medieval astronomers tracked planetary positions with exquisite precision, yet lacked the concept of gravity that would make sense of their observations. Fourier discovered that heat flow could be decomposed into sinusoidal components, but the notion of a function space—where functions themselves become objects of study—lay decades in the future. In each case, implicit understanding preceded explicit conceptualization by generations.
Current AI systems excel at the implicit. A language model trained on mathematical texts develops internal representations that cluster similar concepts, recognize syntactic patterns, and even generate plausible proofs. Yet these representations remain locked in high-dimensional vector spaces, inaccessible to human understanding and incapable of the modular recombination that drives mathematical progress. The challenge is not getting AI to recognize patterns — neural networks do this naturally. The challenge is getting AI to crystallize these patterns into explicit, manipulable concepts.
This challenge of automating the leap from implicit understanding to explicit conceptual compression stands as perhaps the central question for mathematical AI. Without solving it, we risk creating systems that can verify but not discover, that can solve problems but not reshape problem spaces, that can follow mathematics but not lead it into new territories. The question is not whether AI can do mathematics, but whether it can do the kind of mathematics that matters: the kind that doesn’t just solve problems but transforms how we think about them. We now turn to some ideas for achieving this.
From Implicit Patterns to Explicit Concepts
The history of mathematical concepts reveals a striking pattern. Consider how the notion of ”function” emerged from centuries of implicit use. Medieval astronomers computed planetary positions, Renaissance engineers calculated trajectories, and early calculus pioneers found tangent lines using input-output relationships without explicitly naming them. The pattern was everywhere, yet nowhere explicit.
What motivated the transformation of this ubiquitous pattern into the foundational concept we now call ”function”? The answer illuminates how mathematical concepts emerge and suggests how we might teach machines to discover them. We propose that a pattern becomes a concept when making it explicit yields what we might call compositional advantage; when the conceptualization enables efficient compression and generation of mathematical knowledge.
We follow the framework of [EJBL24] and view $Z$ as a corpus of structured data to be understood (current human mathematical output, formal proofs in ZFC, text on the internet, planetary orbits …) and $W$ as an internal, relational language modeling $Z$ through a structure preserving function $f$. We can then formalize the notion of compositional advantage through three critera:
First, the pattern must exhibit high expressive value: $K(Z)$ is large, meaning the mathematical corpus contains rich, complex phenomena worth explaining. The pattern of input-output relationships appeared in astronomy, physics, and pure mathematics, encoding deep regularities across diverse domains.
Second, the pattern must admit high reuse value: $K(Z|W)$ is small, meaning that the complexity of the data $Z$, given the conceptual library $W$ is small. Once ”function” was named, countless phenomena could be redescribed more concisely. The same abstract pattern (mapping inputs to outputs) compressed descriptions across all of mathematics.
Third, the pattern must possess semantic simplicity: $K(f)$ is small, meaning that the complexity of the mapping from $W$ to $Z$ is small. The rules for working with functions ( composition, inversion, restriction ) are remarkably simple yet generate endless complexity.
In this framework, a mathematical concept derives its meaning entirely from its relations to other concepts - how it composes, transforms, and interacts within the mathematical universe. This perspective aligns with category theory’s philosophy: objects gain meaning through relations. Indeed, we might tentatively model mathematical knowledge as a multicategory where concepts are objects and their various relations form different types of morphisms, though this formalization requires further development.
This relational view explains why mathematical concepts exhibit such remarkable portability. When we invoke ’continuity’, we’re not pointing to any specific $\epsilon$-$\delta$ calculation but to a node in a relational network that can be instantiated wherever the proper relations hold; from metric spaces to topologies to categories.
How might AI systems discover such concepts? We propose two complementary approaches that mirror different modes of human mathematical practice.
The first approach involves automated refactoring of existing mathematics. Just as software engineers extract repeated code into reusable functions, an AI system could mine formal mathematical libraries for patterns that appear with systematic variations. When similar proof structures recur across different contexts, the system would suggest extracting a lemma. When multiple theorems share complex hypotheses, it would propose bundling them into a definition. This mirrors how the notion of a ”normal subgroup” might have emerged: mathematicians noticed this property appearing repeatedly in their investigations and crystallized it into a concept.
Concretely, such a system would minimize
$L_{\text{refactor}}=K(Z|W)+\lambda\cdot K(f)$
where the first term encourages high reuse and the second maintains semantic simplicity. The system learns that good concepts are those that appear frequently ($K(Z|W)$ small) while remaining easy to apply ($K(f)$ small).
The second approach pursues exploratory concept discovery, creating novel axiomatic frameworks and evaluating their generative power. Rather than extracting patterns from existing mathematics, the system would propose speculative frameworks $W^{\prime}$ and assess their mathematical value through metrics like unexpected connections: How many previously unrelated areas does $W^{\prime}$ unify?, theorem density: How many non-trivial results emerge from simple axioms? and pattern surprise: Using information theory, how much do discoveries in $W^{\prime}$ violate prior expectations?
This approach formalizes what Grothendieck achieved with schemes, creating an initially alien framework that proved its worth through the mathematics it enabled. The system would optimize
$L_{\text{explore}}=-\frac{K(Z_{\text{discovered}})}{K(Z_{\text{discovered}}|W^{\prime})}+\mu\cdot K(W^{\prime})$
seeking frameworks that enable maximal compression of discovered mathematics while maintaining conceptual parsimony.
Both proposals break from current machine learning paradigms in crucial ways. First, the loss functions involved are computationally expensive (even estimating Kolmogorov complexity is difficult, let alone building an exploratory corpus of $Z_{\text{discovered}}$). Necessarily then, the updates on each loss function have to be significant. Moreover, these proposals aim to rewrite the very language the models output, instead of working within the constraints of a fixed architecture. However, we believe that this difficulty reflects the deeper insight that good concepts are difficult to instantiate and integrate.
The challenge for AI systems is not recognizing patterns, neural networks do this naturally, but crystallizing patterns into explicit, manipulable concepts that enhance mathematical practice. This requires moving beyond prediction accuracy to optimize for conceptual clarity and compositional power. Only then might AI systems not just solve problems but transform problem spaces, not just follow mathematics but lead it into new territories.
The Mathematical Enterprise Reimagined
We began with the Hilbertian program of mechanizing deduction and the resulting transformation of mathematical practice but now face a transformation that may dwarf even this: the emergence of machines that do mathematics as well as, and soon better than, humans. This not the first such transformation but promises to be the most radical yet.
When electronic computers arrived, mathematicians shifted their valuation from computational prowess to abstract proof. When the four-color theorem required computer verification, they debated what constitutes understanding. Each time, the mathematical community redefined what mattered, preserving human relevance by moving the goalposts. The coming AI revolution will demand another such redefinition - but this time, there may be nowhere left to move.
Consider our thought experiment: an AI proves the Riemann Hypothesis with a 10,000-page formal proof that no human can follow. The proof is mechanically verified, each step correct, yet it remains opaque to human understanding. Is this triumph or tragedy? The answer depends on recognizing a fundamental tension in mathematical cognition.
Recall our framework where mathematical knowledge consists of a relational language of concepts $W$ and their semantic interpretation $f:W\to Z$. The total complexity decomposes as:
$K(Z)=K(W)+K(f)+K(Z|W,f)$
Human mathematical practice reflects a severe constraint: our cognitive architecture drastically limits the complexity of $f$ we can execute. With working memory measured in single digits and computational speed in seconds per operation, we cannot mentally perform complex semantic mappings. We compensate by developing rich conceptual vocabularies (large $K(W)$) that enable us to reason with simple mental operations (small $K(f)$).
The epsilon-delta definition exemplifies this tradeoff. Instead of computing limits through infinite processes, we created a concept that reduces verification to local calculations. Category theory pushes this to an extreme: vast conceptual machinery that enables reasoning about complex phenomena through simple diagram chasing. Human mathematics optimizes for $K(f)\ll K(W)$: we prefer elaborate concepts that require minimal mental computation over simple concepts demanding complex calculation.
AI systems face the opposite constraint. With massive parallelism and perfect memory, they can execute semantic mappings of extraordinary complexity. An AI might maintain a sparse conceptual framework (small $K(W)$) while deploying semantic functions of astronomical complexity (large $K(f)$). Where humans prove theorems by discovering conceptual compressions that reduce computational burden, AIs might prove theorems through brute computational force guided by minimal conceptual structure.
This divergence suggests why human-AI mathematical collaboration may prove more difficult than optimists imagine. We’re not simply operating at different speeds but in fundamentally different cognitive regimes.
In the human regime, where $K(f)\ll K(W)$, we develop rich conceptual vocabularies that enable simple mental operations. Our understanding emerges through conceptual manipulation and pattern matching, with proofs valued not merely for their conclusions but for the insight they provide by revealing hidden conceptual structures. We build elaborate theoretical frameworks precisely because they allow us to reason with minimal computational effort.
The AI regime inverts this relationship. With $K(f)\gg K(W)$, these systems can maintain sparse conceptual frameworks while deploying semantic functions of astronomical complexity. Their understanding emerges through massive calculation and systematic search rather than conceptual insight. The proofs they generate may achieve absolute certainty while remaining opaque to human comprehension—correct at every step yet revealing no intelligible structure. The 10,000-page Riemann Hypothesis proof might use familiar number-theoretic concepts but connect them through computational paths no human could trace: each step simple, their composition incomprehensible. The AI hasn’t discovered alien mathematics but rather alien ways of doing familiar mathematics.
Consider Klartag’s recent breakthrough in sphere packing, improving bounds through creative synthesis of stochastic processes and convex geometry. This exemplifies human mathematical virtue: combining distant conceptual frameworks (high $K(W)$) through elegant arguments (low $K(f)$). Yet precisely such synthesis, which humans achieve through conceptual bridges, AIs might accomplish through computational tours de force, checking millions of possibilities until finding connections invisible to conceptual reasoning alone.
This raises profound questions about mathematical understanding. Can we claim to understand a theorem if we grasp the concepts ($W$) but cannot execute the semantic mapping ($f$)? Is there meaningful comprehension without computational capability? The traditional view equates understanding with the ability to reconstruct proofs. But if $K(f)$ exceeds human cognitive capacity, this criterion becomes obsolete.
Three futures seem possible:
First, we might develop conceptual interfaces that translate between cognitive regimes. Just as we created visualizations to understand high-dimensional data, we might create conceptual compressions of AI proofs - not the full argument but a human-comprehensible shadow that captures essential insights. This would require new mathematical practices: extracting conceptual cores from computational arguments, identifying reusable patterns in AI proofs, developing hybrid frameworks where humans provide conceptual guidance while AIs handle computational complexity.
Second, mathematics might bifurcate into human and machine variants. Human mathematics would continue its conceptual tradition, accepting only proofs that provide insight. Machine mathematics would pursue truth through any means, generating a vast corpus of verified but incomprehensible results. The relationship might resemble that between theoretical and experimental physics - distinct but complementary approaches to truth.
Third, and perhaps most likely, we might see a fundamental transformation in what it means to do mathematics. Just as calculators freed us from arithmetic to focus on higher reasoning, AI might free us from proof to focus on something else entirely - perhaps the curation of mathematical aesthetics, the identification of significant questions, or the development of mathematical frameworks that bridge human intuition and machine capability.
The chess analogy, while instructive, ultimately misleads. Chess is a game with fixed rules where humans now play for joy rather than supremacy. Mathematics is not a game but a mode of understanding, one that continuously creates new rules and domains. When machines surpass us, they won’t simply play our game better, they’ll transform the game itself.
Yet perhaps this transformation, like those before it, will reveal new depths in mathematical experience. The undergraduate discovering that continuous functions preserve connectedness will still feel the same ”aha!” moment. The graduate student will still triumph in their first original proof. These experiences persist not because they advance the frontier but because they advance the individual. Mathematics began with humans recognizing patterns in nature and will continue as long as humans find joy in understanding - not all patterns, not the deepest patterns, but our patterns, the ones that fit our finite minds and fill them with wonder.